The most recent settlement T-bond futures price is 110. Which of the following four bonds is...

  1. The most recent settlement T-bond futures price is 110. Which of the following four bonds is cheapest to deliver (use the gross basis)?
  1. Quoted bond price = 123; conversion factor = 1.1000.
  2. Quoted bond price = 155; conversion factor = 1.4000.
  3. Quoted bond price = 145; conversion factor = 1.3000.
  4. Quoted bond price = 137; conversion factor = 1.2200.
Answer #1

Cost of Delivering a Bond = Quoted Bond Price - [Recent Settlement Price * Conversion Factor]

Bond A's Cost of Delivering a Bond = 123 - [110 * 1.1000] = 123 - 121 = 2.00

Bond B's Cost of Delivering a Bond = 155 - [110 * 1.4000] = 155 - 154 = 1.00

Bond C's Cost of Delivering a Bond = 145 - [110 * 1.3000] = 145 - 143 = 2.00

Bond D's Cost of Delivering a Bond = 137 - [110 * 1.2200] = 137 - 134.2 = 2.80

So, the bond in B is therefore cheapest to deliver.